# On Oscillatory Integral Operators Satisfying the cinematic curvature condition

### v1

@@ 3.7. Broad Norm
The broad norm has the following properties; cf. [6, Section 4] for proofs.
@@
The refined decoupling inequality first appeared in [8] as a strengthening of the classical decoupling inequality. In [4, Appendix A], the following refined decoupling inequality was established:
@@ 4. Two-ends reduction
In order to formalize the two-ends and broad structures, we introduce the following notions from [4, Section 6].
@@
[4, Definition 6.1] Let $\Omega$ be the set of horizontal regions introduced above. A shaded incidence triple (or simply triple) $({\bf B},{\bf T};\mathcal{G})$ is the following:
@@
[4, Definition 7.2] Given a triple $({\bf B},{\bf T};\mathcal{G})$, we denote
@@
[4, Definition 6.2] For an $R$-plank $T$, we use $\theta(T)\in\Theta_{R^{-1/2}}$ to denote its directional cap. Let $S$ be a union of $\theta\in\Theta_{R^{-1/2}}$ and let $\bf T$} be a set of $R$-planks. We define
@@
[4, Definition 6.3] Given a triple $({\bf B},{\bf T};\mathcal{G})$, define
@@ 4.1. Algorithm 1
Given $\lambda\lesssim R^{O(1)}$ and $\beta\lesssim K_{\circ}$, define
(4.3) $\mathcal{G}_{\lambda}(T):=\{\omega\in\mathcal{G}(T):\#\{B\in{\bf B}_{1}(T;\mathcal{G}):B\subset\omega\}\sim\lambda\}$
(4.4) ${\bf T}_{\lambda,\beta}=\{T\in{\bf T}:\#\mathcal{G}_{\lambda}(T)\sim\beta\}$
Fix $B\in{\bf B}_{1}$. By pigeonholing, there exists $\lambda(B)$ and $\beta(B)$ such that
$\nu\sim\big\|\sum_{T\in{\bf T}(B;\mathcal{G})}T^{\lambda}f_{T}\big\|_{L^{p}(B)}\leq(\log R)^{O(1)}\big\|\sum_{T\in{\bf T}_{\lambda,\beta}(B;\mathcal{G}_{\lambda(B)})}T^{\lambda}f_{T}\big\|_{L^{p}(B)}$
By dyadic pigeonholing again, we can find $\lambda$, $\beta$, $\nu_{1}$ and ${\bf B}_{2}\subset{\bf B}_{1}$ such that
$\lambda(B)\sim\lambda,\qquad\beta(B)\sim\beta,\qquad\big\|\sum_{T\in{\bf T}_{\lambda,\beta}(B;\mathcal{G}_{\lambda})}T^{\lambda}f_{T}\big\|_{L^{p}(B)}\sim\nu_{1}\qquad\text{for all $B\in{\bf B}_{2}$}$
@@
(4.5) $L^{p}({\bf B}_{1},{\bf T};\mathcal{G};f)\leq(\log R)^{O(1)}L^{p}({\bf B}_{2},{\bf T}_{\lambda,\beta};\mathcal{G}_{\lambda};f)$
@@
$_{\textup{br}}\bigcup_{B\in{\bf B}_{2},B\cap Q\neq\emptyset}{\bf T}_{\lambda,\beta}(B;\mathcal{G}_{\lambda})\sim m\qquad\text{for all $Q\in{\mathcal{Q}}$}$
@@
(4.6) $L^{p}({\bf B}_{2},{\bf T}_{\lambda,\beta};\mathcal{G}_{\lambda};f)\leq(\log R)^{O(1)}L^{p}({\bf B}_{3},{\bf T}_{\lambda,\beta};\mathcal{G}_{\lambda};f)$
@@
$l:=\sup_{T\in{\bf T}_{\lambda,\beta}}\#\{Q\in{\mathcal{Q}}:Q\cap T\neq\emptyset\}$
@@
(4.8) $\mathcal{I}({\bf B}_{1},{\bf T}_{\lambda,\beta};\mathcal{G}_{\lambda})=\sum_{T\in{\bf T}_{\beta}}\#{\bf B}_{1}(T;\mathcal{G}_{\lambda})\lesssim\#{\bf T}_{\lambda,\beta}\cdot\beta\lambda$
@@
(4.9) $L^{p}({\bf B},{\bf T};\mathcal{G};f)\leq(\log R)^{O(1)}L^{p}({\bf B}_{3},{\bf T}_{\lambda,\beta};\mathcal{G}_{\lambda};f)$
@@
[4, Lemma 6.4] Suppose $\beta>100$. Then
${\bf B}_{1}\gtrsim(KK_{\circ})^{-O(1)}lm\lambda$
@@ 4.2. Algorithm 2
Algorithm 2 repeatedly applies Algorithm 1 to establish additional useful quantitative relationships. Let $q:=2+\frac{4}{n-2}$. By a pigeonholing argument and Lemma 3.6, there exists a family ${\mathcal{B}}=\{B\}$ of finitely overlapping $K^{2}$-balls such that
@@
    A set of planks $\mathbb{T}_{\lambda^{(i-1)},\beta^{(i-1)}}^{(i-1)}\subset\mathbb{T}^{(i-1)}$, a shading map $\mathcal{G}_{\lambda^{(i-1)}}^{(i-1)}\subset\mathcal{G}^{(i-1)}$, and a set of $R^{1/2}$-cubes ${\mathcal{Q}}^{(i-1)}$.
We set ${\mathcal{B}}^{(i)}:={\mathcal{B}}_{3}^{(i-1)}$, $\mathbb{T}^{(i)}:=\mathbb{T}_{\lambda^{(i-1)},\beta^{(i-1)}}^{(i-1)}$ and $\mathcal{G}^{(i)}:=\mathcal{G}_{\lambda^{(i-1)}}^{(i-1)}$.
@@
With a slight abuse of notation, we suppress the superscript and denote $\mu^{(N)}$, $\lambda^{(N)}$, $\beta^{(N)}$, $m^{(N)}$, $l^{(N)}$, ${\mathcal{B}}^{(N+1)}$, $\mathbb{T}^{(N+1)}$, $\mathcal{G}^{(N+1)}$ and ${\mathcal{Q}}^{(N)}$ simply by $\mu$, $\lambda$, $\beta$, $m$, $l$, $\mathcal{B}$}, $\mathbb{T}$, $\mathcal{G}$ and $\mathcal{Q}$}.
@@ 5. Complete the proof
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1), (3.2), (3.22) and (3.23), then for all $p$ satisfies (1.2), the estimate
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}\lesssim_{\varepsilon,p,\phi}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{p}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{p}}$
@@
We work under the same setting as in Section 4. Cover $B_{R}^{n}(0)$ by finite-overlapping $R/K_{\circ}$-balls $B_{k}$. For each $B_{k}$ define $\mathbb{T}_{k}$ by
@@ 5.1. Case: One-end
(5.2) $\|T^{\lambda}f_{k}\|_{L^{p}(B_{k})}^{p}\leq C_{\varepsilon}(\frac{R}{K_{\circ}})^{p\varepsilon}\|f_{k}\|_{2}^{2}\,\max_{\theta\in\Theta_{K_{\circ}^{1/2}R^{-1/2}}}\|f_{k}\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}+{\rm RapDec}(R)\,\|f\|_{2}^{p}\\ \leq C_{\varepsilon}(\frac{R}{K_{\circ}})^{p\varepsilon}\|f_{k}\|_{2}^{2}\,\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}+{\rm RapDec}(R)\,\|f\|_{2}^{p}$
@@
(5.3) $\max_{\theta\in\Theta_{R^{-1/2}}}\|f_{k}\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}\lesssim\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}+{\rm RapDec}(R)\,\|f\|_{2}^{p-2}$
@@
$\|T^{\lambda}f_{k}\|_{L^{p}(B_{k})}^{p}\leq C_{\varepsilon}(\frac{R}{K_{\circ}})^{p\varepsilon}\|f_{k}\|_{2}^{2}\,\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}+{\rm RapDec}(R)\,\|f\|_{2}^{p}$
@@
(5.4) $\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim C_{\varepsilon}K_{\circ}^{-p\varepsilon}(\log R)^{O_{\varepsilon}(1)}R^{p\varepsilon}\sum_{B_{k}}\|f_{k}\|_{2}^{2}\,\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}\\ +{\rm RapDec}(R)\,\|f\|_{2}^{p}$
Since $\beta\leq 100$, each wave packet $f_{T}$ appears in the sum defining $O(1)$ different functions $f_{k}$. Consequently, Lemma 3.7 yields
$\sum_{k}\|f_{k}\|_{2}^{2}\lesssim\|f\|_{2}^{2}$
@@
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim C_{\varepsilon}K_{\circ}^{-p\varepsilon}(\log R)^{O_{\varepsilon}(1)}R^{p\varepsilon}\|f\|_{2}^{2}\,\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-2}\\ +{\rm RapDec}(R)\,\|f\|_{2}^{p}$
@@ 5.2. Two-ends: Reduction to Broad Estimate
$\min_{v_{1},\dots,v_{A}\in S^{n-1}}\max_{\begin{subarray}{c}\tau\in\Theta_{K^{-1}}\\ \measuredangle(G_{0}(y/\lambda;\omega_{\tau},1),v_{i})\gtrsim K^{-2}\text{for all $1\leq i\leq A$}\end{subarray}}\int_{B_{K^{2}}^{n}(y)}|T^{\lambda}f_{k,\tau}|^{p}\\ \text{where}\quad f_{k,\tau}:=(f_{k})_{\tau}$
@@
$\|T^{\lambda}f_{k}\|_{L^{p}(B_{K^{2}}^{n}(y))}^{p}\lesssim\sum_{1\leq i\leq A}\|T^{\lambda}(\sum_{{\begin{subarray}{c}\tau\in\Theta_{K^{-1}}\\ \measuredangle(G_{0}(y/\lambda;\omega_{\tau},1),v_{i})\lesssim K^{-2}\end{subarray}}}f_{k,\tau})\|_{L^{p}(B_{K^{2}}^{n}(y))}^{p}\\ +K^{O(1)}\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(B_{K^{2}}^{n}(y))}^{p}$
For each $i$, there are $O(1)$ $K^{-1}$-caps satisfies $\measuredangle(G_{0}(y/\lambda;\omega_{\tau},1),v_{i})\lesssim K^{-2}$, so
@@
$\|T^{\lambda}f_{k,\tau}\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim_{\phi}K^{\,2-(n-2)(p-2)-p\varepsilon}R^{p\varepsilon}\,\|f_{k,\tau}\|_{2}^{2}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,p-2}\\ +{\rm RapDec}(R)\,\|f\|_{2}^{p}$
@@
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\lessapprox_{\phi}K_{\circ}^{O(1)}K^{\,2-(n-2)(p-2)-p\varepsilon}R^{p\varepsilon}\,\|f\|_{2}^{2}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,p-2}\\ +\big(K_{\circ}K\big)^{O(1)}\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(X)}^{p}+{\rm RapDec}(R)\,\|f\|_{2}^{p}\vskip 12.0pt plus 4.0pt minus 4.0pt$
If $p>2+\frac{2}{n-2}$, then the narrow term vanishes, and hence it remains to show:
(5.7) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(X)}^{p}\lesssim_{\phi,\varepsilon}R^{\varepsilon}\|f\|_{2}^{2}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,p-2}$

### v2

@@ 3.7. Broad Norm
The broad norm has the following properties; cf. [7, Section 4] for proofs.
@@
The refined decoupling inequality first appeared in [9] as a strengthening of the classical decoupling inequality. In [5, Appendix A], the following refined decoupling inequality was established:
@@ 4. Two-ends reduction
In order to formalize the two-ends and broad structures, we introduce the following notions from [5, Section 6].
@@
[5, Definition 6.1] Let $\Omega$ be the set of horizontal regions introduced above. A shaded incidence triple (or simply triple) $({\bf B},{\bf T};\mathcal{G})$ is the following:
@@
[5, Definition 7.2] Given a triple $({\bf B},{\bf T};\mathcal{G})$, we denote
@@
[5, Definition 6.2] For an $R$-plank $T$, we use $\theta(T)\in\Theta_{R^{-1/2}}$ to denote its directional cap. Let $S$ be a union of $\theta\in\Theta_{R^{-1/2}}$ and let $\bf T$} be a set of $R$-planks. We define
@@
[5, Definition 6.3] Given a triple $({\bf B},{\bf T};\mathcal{G})$, define
@@ 4.1. Algorithm 1
Given $W\lesssim R^{O(1)}$ and $\beta\lesssim K_{\circ}$, define
(4.3) $\mathcal{G}_{W}(T):=\{\omega\in\mathcal{G}(T):\#\{B\in{\bf B}_{1}(T;\mathcal{G}):B\subset\omega\}\sim W\}$
(4.4) ${\bf T}_{W,\beta}=\{T\in{\bf T}:\#\mathcal{G}_{W}(T)\sim\beta\}$
Fix $B\in{\bf B}_{1}$. By pigeonholing, there exists $W(B)$ and $\beta(B)$ such that
$\nu\sim\big\|\sum_{T\in{\bf T}(B;\mathcal{G})}T^{\lambda}f_{T}\big\|_{L^{p}(B)}\leq(\log R)^{O(1)}\big\|\sum_{T\in{\bf T}_{W(B),\beta}(B;\mathcal{G}_{W(B)})}T^{\lambda}f_{T}\big\|_{L^{p}(B)}$
By dyadic pigeonholing again, we can find $W$, $\beta$, $\nu_{1}$ and ${\bf B}_{2}\subset{\bf B}_{1}$ such that
$W(B)\sim W,\qquad\beta(B)\sim\beta,\qquad\big\|\sum_{T\in{\bf T}_{W,\beta}(B;\mathcal{G}_{W})}T^{\lambda}f_{T}\big\|_{L^{p}(B)}\sim\nu_{1}\qquad\text{for all $B\in{\bf B}_{2}$}$
@@
(4.5) $L^{p}({\bf B}_{1},{\bf T};\mathcal{G};f)\leq(\log R)^{O(1)}L^{p}({\bf B}_{2},{\bf T}_{W,\beta};\mathcal{G}_{W};f)$
@@
$_{\textup{br}}\bigcup_{B\in{\bf B}_{2},B\cap Q\neq\emptyset}{\bf T}_{W,\beta}(B;\mathcal{G}_{W})\sim m\qquad\text{for all $Q\in{\mathcal{Q}}$}$
@@
(4.6) $L^{p}({\bf B}_{2},{\bf T}_{W,\beta};\mathcal{G}_{W};f)\leq(\log R)^{O(1)}L^{p}({\bf B}_{3},{\bf T}_{W,\beta};\mathcal{G}_{W};f)$
@@
$l:=\sup_{T\in{\bf T}_{W,\beta}}\#\{Q\in{\mathcal{Q}}:Q\cap T\neq\emptyset\}$
@@
(4.8) $\mathcal{I}({\bf B}_{1},{\bf T}_{W,\beta};\mathcal{G}_{W})=\sum_{T\in{\bf T}_{\beta}}\#{\bf B}_{1}(T;\mathcal{G}_{W})\lesssim\#{\bf T}_{W,\beta}\cdot\beta W$
@@
(4.9) $L^{p}({\bf B},{\bf T};\mathcal{G};f)\leq(\log R)^{O(1)}L^{p}({\bf B}_{3},{\bf T}_{W,\beta};\mathcal{G}_{W};f)$
@@
[5, Lemma 6.4] Suppose $\beta>100$. Then
${\bf B}_{1}\gtrsim(KK_{\circ})^{-O(1)}Wlm$
@@ 4.2. Algorithm 2
Algorithm 2 repeatedly applies Algorithm 1 to establish additional useful quantitative relationships. By a pigeonholing argument and Lemma 3.6, there exists a family ${\mathcal{B}}=\{B\}$ of finitely overlapping $K^{2}$-balls such that
@@
    A set of planks $\mathbb{T}_{W^{(i-1)},\beta^{(i-1)}}^{(i-1)}\subset\mathbb{T}^{(i-1)}$, a shading map $\mathcal{G}_{W^{(i-1)}}^{(i-1)}\subset\mathcal{G}^{(i-1)}$, and a set of $R^{1/2}$-cubes ${\mathcal{Q}}^{(i-1)}$.
We set ${\mathcal{B}}^{(i)}:={\mathcal{B}}_{3}^{(i-1)}$, $\mathbb{T}^{(i)}:=\mathbb{T}_{W^{(i-1)},\beta^{(i-1)}}^{(i-1)}$ and $\mathcal{G}^{(i)}:=\mathcal{G}_{W^{(i-1)}}^{(i-1)}$.
@@
With a slight abuse of notation, we suppress the superscript and denote $\mu^{(N)}$, $W^{(N)}$, $\beta^{(N)}$, $m^{(N)}$, $l^{(N)}$, ${\mathcal{B}}^{(N+1)}$, $\mathbb{T}^{(N+1)}$, $\mathcal{G}^{(N+1)}$ and ${\mathcal{Q}}^{(N)}$ simply by $\mu$, $W$, $\beta$, $m$, $l$, $\mathcal{B}$}, $\mathbb{T}$, $\mathcal{G}$ and $\mathcal{Q}$}.
@@ 5. Complete the proof
Suppose $T^{\lambda}$ is an oscillatory integral operator satisfying the cinematic curvature condition of type $\mathbf{1}$. Assume (3.1), (3.2), (3.22) and (3.23), then for all $p$ satisfies (1.4) and $r$ satisfies (1.3), the estimate
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}\lesssim_{\varepsilon,p,\phi}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{r}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{r}}$
@@
We work under the same setting as in Section 4. By interpolating with $L^{1}\mapsto L^{\infty}$ bound and and applying Hölder’s inequality, we may assume that
(5.1) $p=2+\frac{8}{3n-5},\qquad\frac{r}{r-1}=\frac{n-2}{n}p$
Cover $B_{R}^{n}(0)$ by finite-overlapping $R/K_{\circ}$-balls $B_{k}$. For each $B_{k}$ define $\mathbb{T}_{k}$ by
@@ 5.1. Case: One-end
(5.3) $\|T^{\lambda}f_{k}\|_{L^{p}(B_{k})}^{p}\leq C_{\varepsilon}(\frac{R}{K_{\circ}})^{p\varepsilon}\|f_{k}\|_{2}^{\frac{2p}{r}}\,\max_{\theta\in\Theta_{K_{\circ}^{1/2}R^{-1/2}}}\|f_{k}\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-\frac{2p}{r}}+{\rm RapDec}(R)\,\|f\|_{2}^{p}$
@@
(5.4) $\max_{\theta\in\Theta_{R^{-1/2}}}\|f_{k}\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-\frac{2p}{r}}\lesssim\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-\frac{2p}{r}}+{\rm RapDec}(R)\,\|f\|_{2}^{p-\frac{2p}{r}}$
@@
$\|T^{\lambda}f_{k}\|_{L^{p}(B_{k})}^{p}\leq C_{\varepsilon}(\frac{R}{K_{\circ}})^{p\varepsilon}\|f_{k}\|_{2}^{\frac{2p}{r}}\,\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-\frac{2p}{r}}+{\rm RapDec}(R)\,\|f\|_{2}^{p}$
@@
(5.5) $\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim C_{\varepsilon}K_{\circ}^{-p\varepsilon}(\log R)^{O_{\varepsilon}(1)}R^{p\varepsilon}\sum_{B_{k}}\|f_{k}\|_{2}^{\frac{2p}{r}}\,\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-\frac{2p}{r}}\\ +{\rm RapDec}(R)\,\|f\|_{2}^{p}$
Since $\beta\leq 100$, each wave packet $f_{T}$ appears in the sum defining $O(1)$ different functions $f_{k}$. Also, from (5.1) we see that $r\leq p$. Consequently, Lemma 3.7 yields
$\sum_{k}\|f_{k}\|_{2}^{\frac{2p}{r}}\lesssim\|f\|_{2}^{\frac{2p}{r}}$
@@
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim C_{\varepsilon}K_{\circ}^{-p\varepsilon}(\log R)^{O_{\varepsilon}(1)}R^{p\varepsilon}\|f\|_{2}^{\frac{2p}{r}}\,\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-\frac{2p}{r}}\\ +{\rm RapDec}(R)\,\|f\|_{2}^{p}$
@@ 5.2. Two-ends: Reduction to Broad Estimate
$\min_{v_{1},\dots,v_{A}\in S^{n-1}}\max_{\begin{subarray}{c}\tau\in\Theta_{K^{-1}}\\ \measuredangle(G_{0}(y/\lambda;\omega_{\tau},1),v_{i})\gtrsim K^{-1}\text{for all $1\leq i\leq A$}\end{subarray}}\int_{B_{K^{2}}^{n}(y)}|T^{\lambda}f_{k,\tau}|^{p}\\ \text{where}\quad f_{k,\tau}:=(f_{k})_{\tau}$
@@
$\|T^{\lambda}f_{k}\|_{L^{p}(B_{K^{2}}^{n}(y))}^{p}\lesssim\sum_{1\leq i\leq A}\|T^{\lambda}(\sum_{{\begin{subarray}{c}\tau\in\Theta_{K^{-1}}\\ \measuredangle(G_{0}(y/\lambda;\omega_{\tau},1),v_{i})\lesssim K^{-1}\end{subarray}}}f_{k,\tau})\|_{L^{p}(B_{K^{2}}^{n}(y))}^{p}\\ +K^{O(1)}\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(B_{K^{2}}^{n}(y))}^{p}$
For each $i$, there are $O(1)$ $K^{-1}$-caps satisfies $\measuredangle(G_{0}(y/\lambda;\omega_{\tau},1),v_{i})\lesssim K^{-1}$, so
@@
$\|T^{\lambda}f_{k,\tau}\|_{L^{p}(B_{R}^{n}(0))}^{p}\lesssim_{\phi}K^{n-(n-2)(p-\frac{p}{r})}R^{p\varepsilon}\,\|f_{k,\tau}\|_{2}^{\frac{2p}{r}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-\frac{2p}{r}}\\ +{\rm RapDec}(R)\,\|f\|_{2}^{p}$
@@
$\|T^{\lambda}f\|_{L^{p}(B_{R}^{n}(0))}^{p}\lessapprox_{\phi}K_{\circ}^{O(1)}K^{n-(n-2)(p-\frac{p}{r})-p\varepsilon}R^{p\varepsilon}\,\|f\|_{2}^{\frac{2p}{r}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{p-\frac{2p}{r}}\\ +\big(K_{\circ}K\big)^{O(1)}\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(X)}^{p}+{\rm RapDec}(R)\,\|f\|_{2}^{p}\vskip 12.0pt plus 4.0pt minus 4.0pt$
Under (5.1) the narrow term vanishes, and thus it remains to show:
(5.8) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(X)}\lesssim_{\phi,\varepsilon}R^{\varepsilon}\|f\|_{2}^{\frac{2}{r}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{r}}$